# Xiaoyue Wang  Jianyou Wang11footnotemark: 1  Weili Cao11footnotemark: 1  Kaicheng Wang11footnotemark: 1 Ramamohan Paturi  Leon Bergen22footnotemark: 2

### v1

@@ Appendix A Experiment Details
Score+OSH-7b 67.5 $\pm$ 1.4 9.5 $\pm$ 1.4 20.6 $\pm$ 2.5 24.9 $\pm$ 4.1 43.0 $\pm$ 4.0 50.8 $\pm$ 4.9 32.5 $\pm$ 2.3 10.7 $\pm$ 1.3 20.7 $\pm$ 3.0 25.0 $\pm$ 4.4
36.9
Reason+OSH-7b 65.9 $\pm$ 1.1 10.9 $\pm$ 1.4 26.3 $\pm$ 3.4 32.0 $\pm$ 4.9 22.2 $\pm$ 3.6 22.1 $\pm$ 4.3 27.5 $\pm$ 2.6 11.0 $\pm$ 1.8 12.8 $\pm$ 2.5 11.8 $\pm$ 3.2
30.9
Subtask+OSH-7b 74.8 $\pm$ 1.3 15.5 $\pm$ 2.3 14.1 $\pm$ 2.1 20.9 $\pm$ 4.0 44.4 $\pm$ 3.7 49.0 $\pm$ 4.9 32.4 $\pm$ 2.5 11.0 $\pm$ 1.3 24.7 $\pm$ 3.1 29.1 $\pm$ 4.5
38.1
Score+OLM2-7b 71.6 $\pm$ 1.4 14.1 $\pm$ 1.9 27.0 $\pm$ 3.2 40.1 $\pm$ 4.9 30.5 $\pm$ 3.8 34.7 $\pm$ 4.8 29.7 $\pm$ 2.8 10.1 $\pm$ 1.5 10.9 $\pm$ 2.4 13.9 $\pm$ 3.3
33.9
Reason+OLM2-7b 72.1 $\pm$ 1.2 13.8 $\pm$ 1.7 21.3 $\pm$ 3.3 27.8 $\pm$ 4.6 21.4 $\pm$ 3.1 28.1 $\pm$ 4.4 31.0 $\pm$ 2.5 9.9 $\pm$ 1.4 11.5 $\pm$ 2.3 12.0 $\pm$ 3.2
31.5
Subtask+OLM2-7b 69.2 $\pm$ 1.4 11.8 $\pm$ 1.7 16.4 $\pm$ 2.4 18.8 $\pm$ 3.8 29.2 $\pm$ 3.4 32.9 $\pm$ 4.4 25.1 $\pm$ 2.2 8.6 $\pm$ 1.3 10.5 $\pm$ 2.6 11.8 $\pm$ 3.1
30.1
Score+OLM2-13b 77.3 $\pm$ 0.9 14.5 $\pm$ 1.6 15.4 $\pm$ 2.4 21.9 $\pm$ 4.0 28.1 $\pm$ 3.3 31.9 $\pm$ 4.5 30.0 $\pm$ 2.4 9.9 $\pm$ 1.1 16.8 $\pm$ 2.5 18.0 $\pm$ 3.7
33.5
Reason+OLM2-13b 60.1 $\pm$ 1.4 7.5 $\pm$ 1.4 31.4 $\pm$ 3.0 43.1 $\pm$ 4.9 18.7 $\pm$ 2.4 27.0 $\pm$ 4.4 24.2 $\pm$ 2.5 7.7 $\pm$ 1.1 9.0 $\pm$ 1.8 8.9 $\pm$ 2.8
28.7
Subtask+OLM2-13b 71.3 $\pm$ 1.3 14.1 $\pm$ 1.7 12.4 $\pm$ 2.3 14.9 $\pm$ 3.5 48.7 $\pm$ 3.9 60.9 $\pm$ 5.0 29.4 $\pm$ 2.3 10.7 $\pm$ 1.3 17.6 $\pm$ 2.7 21.0 $\pm$ 4.0
35.9
Score+OLM2-70b 75.1 $\pm$ 1.2 15.1 $\pm$ 2.0 40.3 $\pm$ 3.4 49.8 $\pm$ 5.0 60.8 $\pm$ 3.7 70.0 $\pm$ 4.7 34.9 $\pm$ 2.6 12.1 $\pm$ 1.6 12.2 $\pm$ 2.6 14.9 $\pm$ 3.4
44.7
Reason+OLM2-70b 70.1 $\pm$ 1.4 14.5 $\pm$ 1.9 42.8 $\pm$ 3.2 55.0 $\pm$ 4.9 43.4 $\pm$ 4.1 48.0 $\pm$ 5.0 31.3 $\pm$ 2.2 9.9 $\pm$ 1.2 12.7 $\pm$ 2.6 15.9 $\pm$ 3.5
40.1
Subtask+OLM2-70b 76.7 $\pm$ 1.2 16.3 $\pm$ 1.9 19.3 $\pm$ 2.7 27.0 $\pm$ 4.4 49.1 $\pm$ 3.8 60.1 $\pm$ 4.8 32.8 $\pm$ 2.9 10.5 $\pm$ 1.3 19.6 $\pm$ 3.2 21.0 $\pm$ 4.0
39.5
Score+OGPT4 79.9 $\pm$ 1.2 19.3 $\pm$ 2.0 51.6 $\pm$ 2.7 70.0 $\pm$ 4.6 83.3 $\pm$ 3.1 90.9 $\pm$ 2.8 43.4 $\pm$ 2.4 17.2 $\pm$ 1.6 54.1 $\pm$ 3.3 70.1 $\pm$ 4.6
62.5
Reason+OGPT4 74.9 $\pm$ 1.3 18.1 $\pm$ 2.3 59.9 $\pm$ 2.9 76.8 $\pm$ 4.2 74.9 $\pm$ 3.6 82.8 $\pm$ 3.7 45.7 $\pm$ 3.1 17.0 $\pm$ 1.7 39.9 $\pm$ 3.3 51.0 $\pm$ 5.0
59.1
Subtask+OGPT4 78.5 $\pm$ 1.2 18.5 $\pm$ 1.9 69.5 $\pm$ 3.0 85.9 $\pm$ 3.5 79.6 $\pm$ 3.5 85.9 $\pm$ 3.3 43.5 $\pm$ 2.9 16.5 $\pm$ 1.6 53.2 $\pm$ 3.5 62.9 $\pm$ 4.8 64.9
Table 7: nDCG@10 and Recall@5 for the benchmark datasets. Bold indicates $p>0.05$ compared to the highest numerical value indicated in red. Error bars are standard errors. The notation $+$O indicates task objective awareness. LM2 stands for Llama2 Touvron et al. (2023). SH stands for StripedHyena Poli et al. (2023)
@@
All embedding models have 350M parameters or less. Promptagator models are E5-v2-L checkpoint finetuned on promptagator-styled synthetic data, introduced by Dai et al. (2022). No data for Clinical-Trial and RELIC is available for Promptagator models. Other embedding models are not fine-tuned for IR tasks.
@@ A.1 Baselines
For pretrained embedding models, we choose several recent state-of-the-art models as well as several that have been extensively benchmarked on other tasks: E5-Large-v2 Wang et al. (2022), SimCSE-Large Gao et al. (2021), SPECTER-v2 Singh et al. (2022), ROBERTA-Large Liu et al. (2019), SPLADE-v2 Formal et al. (2021), and SPLADE++ Formal et al. (2022). There are several encoder-decoder models specifically trained for information retrieval: monoT5-3B Nogueira et al. (2020), which is supervised on MS MARCO, and TART-1.5B Asai et al. (2023), which is instruction-tuned and can incorporate task descriptions. For DORIS-MAE, ArguAna, and WhatsThatBook, we also report results for E5-Large-v2 fine-tuned using synthetic data generated by Promptagator Dai et al. (2022). Additionally, we evaluate a decoder-only model RankLLaMA-7B Ma et al. (2023a) trained on MS MARCO. We also evaluate two other generative models: Llama2 chat Touvron et al. (2023), which is one of the best open-sourced LLMs, and StripedHyena-Nous-7B chat Poli et al. (2023), which is a competitive alternative LLM with a hybrid architecture between multi-headed attention and gated convolution.
@@ Appendix B Task Objectives for BIRCO
BIRCO DORIS-MAE 18.2 100 Multi-level 14.7 42.2 72.0 14.3 40.5
ArguAna 1.0 100 Binary 62.0 96.0 43.6 33.2 34.7
WhatsThatBook 1.0 100 Binary 40.0 64.0 36.8 32.0 33.8
Clinical-Trial 20.9 100 3-level 10.6 37.8 29.4 34.4 37.4
RELIC 1.0 100 Binary 15.0 43.0 11.2 8.9 12.2
Average 8.4 100 - 28.5 56.6 38.6 24.6 31.7
@@ Appendix C The Effects of Candidate Pool Construction
Clinical-Trial 29.4 $\pm$ 2.7 10.5 $\pm$ 1.7 32.4 $\pm$ 2.6 13.1 $\pm$ 1.9
RELIC 11.1 $\pm$ 2.6 14.9 $\pm$ 3.4 8.2 $\pm$ 0.4 10.1 $\pm$ 0.5

### v2

@@ Appendix A Experiment Details
Score+OSH 67.5 $\pm$ 1.4 9.5 $\pm$ 1.4 21.6 $\pm$ 2.6 23.1 $\pm$ 4.2 43.0 $\pm$ 4.0 50.8 $\pm$ 4.9 31.8 $\pm$ 2.7 10.8 $\pm$ 1.5 21.1 $\pm$ 3.0 24.8 $\pm$ 4.4 37.0
Reason+OSH 65.9 $\pm$ 1.1 10.9 $\pm$ 1.4 27.7 $\pm$ 3.5 32.0 $\pm$ 4.6 23.1 $\pm$ 3.7 23.1 $\pm$ 4.4 29.2 $\pm$ 2.8 13.8 $\pm$ 2.5 12.7 $\pm$ 2.4 11.9 $\pm$ 3.2 31.7
Subtask+OSH 74.8 $\pm$ 1.3 15.5 $\pm$ 2.3 16.3 $\pm$ 2.2 20.0 $\pm$ 4.0 43.5 $\pm$ 3.7 53.0 $\pm$ 5.0 32.3 $\pm$ 2.9 10.9 $\pm$ 1.7 24.7 $\pm$ 3.2 29.1 $\pm$ 4.4 38.3
Score+OLM2-7b 71.6 $\pm$ 1.4 14.1 $\pm$ 1.9 29.3 $\pm$ 3.2 40.1 $\pm$ 4.8 30.5 $\pm$ 3.8 34.7 $\pm$ 4.8 31.1 $\pm$ 2.7 11.2 $\pm$ 1.6 10.9 $\pm$ 2.4 14.0 $\pm$ 3.5 34.7
Reason+OLM2-7b 72.1 $\pm$ 1.2 13.8 $\pm$ 1.7 21.9 $\pm$ 3.4 27.0 $\pm$ 4.5 21.4 $\pm$ 3.1 28.1 $\pm$ 4.4 29.9 $\pm$ 2.7 9.7 $\pm$ 1.5 10.1 $\pm$ 2.2 10.0 $\pm$ 3.1 31.1
Subtask+OLM2-7b 69.2 $\pm$ 1.4 11.8 $\pm$ 1.7 22.3 $\pm$ 2.5 27.9 $\pm$ 4.5 29.2 $\pm$ 3.5 31.1 $\pm$ 4.5 28.7 $\pm$ 2.4 10.0 $\pm$ 1.2 9.4 $\pm$ 2.2 11.9 $\pm$ 3.2 31.8
Score+OLM2-13b 77.3 $\pm$ 0.9 14.5 $\pm$ 1.6 15.0 $\pm$ 2.2 24.1 $\pm$ 4.2 29.6 $\pm$ 3.4 34.0 $\pm$ 4.6 30.6 $\pm$ 2.6 10.5 $\pm$ 1.3 16.2 $\pm$ 2.6 18.0 $\pm$ 3.9 33.7
Reason+OLM2-13b 60.1 $\pm$ 1.4 7.5 $\pm$ 1.4 32.6 $\pm$ 3.1 45.1 $\pm$ 5.0 19.4 $\pm$ 2.4 28.0 $\pm$ 4.5 25.4 $\pm$ 2.1 8.7 $\pm$ 1.2 9.1 $\pm$ 1.9 8.9 $\pm$ 2.7 29.3
Subtask+OLM2-13b 71.3 $\pm$ 1.3 14.1 $\pm$ 1.7 18.1 $\pm$ 2.7 20.8 $\pm$ 4.1 52.8 $\pm$ 3.9 61.9 $\pm$ 5.0 27.7 $\pm$ 2.5 10.7 $\pm$ 1.4 19.7 $\pm$ 2.8 20.0 $\pm$ 4.0 37.9
Score+OLM2-70b 75.1 $\pm$ 1.2 15.1 $\pm$ 2.0 41.6 $\pm$ 3.4 53.0 $\pm$ 4.9 60.8 $\pm$ 3.7 70.0 $\pm$ 4.7 36.3 $\pm$ 2.8 12.6 $\pm$ 1.7 11.7 $\pm$ 2.5 14.9 $\pm$ 3.6 45.1
Reason+OLM2-70b 70.1 $\pm$ 1.4 14.5 $\pm$ 1.9 45.0 $\pm$ 3.3 60.1 $\pm$ 4.8 43.4 $\pm$ 4.1 48.0 $\pm$ 5.0 32.6 $\pm$ 2.5 11.6 $\pm$ 1.5 12.1 $\pm$ 2.5 15.9 $\pm$ 3.7 40.6
Subtask+OLM2-70b 76.7 $\pm$ 1.2 16.3 $\pm$ 1.9 21.0 $\pm$ 2.9 25.0 $\pm$ 4.4 42.6 $\pm$ 3.7 50.2 $\pm$ 5.0 33.3 $\pm$ 3.0 11.8 $\pm$ 1.8 17.3 $\pm$ 2.8 21.0 $\pm$ 4.0 38.2
Score+OGPT4 79.9 $\pm$ 1.2 19.3 $\pm$ 2.0 53.2 $\pm$ 2.8 70.8 $\pm$ 4.6 83.3 $\pm$ 3.1 90.9 $\pm$ 2.8 43.7 $\pm$ 2.6 19.6 $\pm$ 2.5 56.9 $\pm$ 3.2 73.2 $\pm$ 4.3 63.4
Reason+OGPT4 74.9 $\pm$ 1.3 18.3 $\pm$ 2.3 61.5 $\pm$ 3.0 76.1 $\pm$ 4.4 75.2 $\pm$ 3.6 82.8 $\pm$ 3.7 44.6 $\pm$ 3.1 16.4 $\pm$ 1.9 41.7 $\pm$ 3.3 52.8 $\pm$ 4.9 59.6
Subtask+OGPT4 78.5 $\pm$ 1.2 18.5 $\pm$ 1.9 70.8 $\pm$ 3.1 83.9 $\pm$ 3.6 77.9 $\pm$ 3.5 85.9 $\pm$ 3.4 44.3 $\pm$ 2.9 19.5 $\pm$ 2.5 53.3 $\pm$ 3.7 66.0 $\pm$ 4.6 65.0
Table 7: nDCG@10 and Recall@5 for the benchmark datasets. Error bars are standard errors. The notation $+$O indicates task objective awareness. LM2 stands for Llama2 Touvron et al. (2023). SH stands for StripedHyena Poli et al. (2023)
@@
All embedding models have 350M parameters or less. Promptagator models are E5-v2-L checkpoints finetuned on promptagator-styled Dai et al. (2022) synthetic data, introduced by Wang et al. (2024). No data for Clinical-Trial and RELIC is available for Promptagator models. Other embedding models are not fine-tuned for IR tasks.
@@ A.1 Baselines
For pretrained embedding models, we choose several recent state-of-the-art models as well as several that have been extensively benchmarked on other tasks: E5-Large-v2 Wang et al. (2022), SimCSE-Large Gao et al. (2021), SPECTER-v2 Singh et al. (2022), ROBERTA-Large Liu et al. (2019), SPLADE-v2 Formal et al. (2021), and SPLADE++ Formal et al. (2022). There are several encoder-decoder models specifically trained for information retrieval: monoT5-3B Nogueira et al. (2020), which is supervised on MS MARCO, and TART-1.5B Asai et al. (2023), which is instruction-tuned and can incorporate task descriptions. For DORIS-MAE, ArguAna, and WhatsThatBook, we also report results for E5-Large-v2 checkpoints fine-tuned using synthetic data generated by Promptagator Dai et al. (2022) from Wang et al. (2024). Additionally, we evaluate a decoder-only model RankLLaMA-7B Ma et al. (2023a) trained on MS MARCO. We also evaluate two other generative models: Llama2 chat Touvron et al. (2023), which is one of the best open-sourced LLMs, and StripedHyena-Nous-7B chat Poli et al. (2023), which is a competitive alternative LLM with a hybrid architecture between multi-headed attention and gated convolution.
@@ Appendix B Task Objectives for BIRCO
BIRCO DORIS-MAE 18.2 110.6 Multi-level 14.7 42.2 72.0 14.3 40.5
ArguAna 1.0 50.0 Binary 62.0 96.0 43.6 33.2 34.7
WhatsThatBook 1.0 50.4 Binary 40.0 64.0 36.8 32.0 33.8
Clinical-Trial 20.9 68.4 3-level 10.6 37.8 29.4 34.4 37.4
RELIC 1.0 50.6 Binary 15.0 43.0 11.2 8.9 12.2
Average 8.4 66.0 - 28.5 56.6 38.6 24.6 31.7
@@ Appendix C The Effects of Candidate Pool Construction
Clinical-Trial 30.4 $\pm$ 2.6 11.3 $\pm$ 1.8 32.4 $\pm$ 2.6 13.1 $\pm$ 1.9
RELIC 10.9 $\pm$ 2.5 14.9 $\pm$ 3.6 8.2 $\pm$ 0.4 10.1 $\pm$ 0.5